### v1

# A note on best $n$-term approximation for generalized Wiener classes

Abstract
We determine the best $n$-term approximation of generalized Wiener model classes in a Hilbert space $H$. This theory is then applied to several special cases.
1 Introduction
[…]
$f=\sum_{j=1}^{\infty}f_{j}\phi_{j},\quad\hbox{where}\quad\|f\|^{2}=\sum_{j=1}^{\infty}|f_{j}|^{2}.$
We are concerned with $n$ term approximation of the elements $f\in H$. We denote by $\Sigma_{n}:=\{S=\sum_{j\in\Lambda}c_{j}e_{j}:\,\Lambda\subset\mathbb{N},\,|\Lambda|=n\}$ and let
$\sigma_{n}(f):=\inf_{S\in\Sigma_{n}}\|f-S\|,\quad n\geq 1,$
be the error of $n$-term approximation of $f$. Given any $f\in H$, a best $n$-term approximation $S_{n}$ of $f$ is given by
$S_{n}=S_{n}(f):=\sum_{j\in\Lambda_{n}}f_{j}\phi_{j},$
where $\Lambda_{n}:=\Lambda_{n}(f)$ is a set of $n$ indices $j$ for which $|f_{j}|\geq|f_{i}|$ whenever $j\in\Lambda$ and $i\notin\Lambda$. Even though the set $\Lambda_{n}(f)$ is not uniquely defined because of possible ties in terms of the size of the absolute values of the coefficients $f_{j}$, the error of approximation $\sigma_{n}(f)$ is uniquely defined.
We are interested in model classes $K\subset H$ that are given by imposing conditions on the coefficients $(f_{j})$ of $f$. For such sets $K$, we define
$\sigma_{n}(K):=\sup_{f\in K}\sigma_{n}(f)$ (1.1)
and we are interested in the asymptotic decay of $\sigma_{n}(K)\to 0$ as $n\to\infty$.
Since we are only considering the approximation to take place in $H$, in going further it is sufficient to consider only the case
$H=\ell_{2}:=\left\{{\bf x}=(x_{1},x_{2},\ldots,):\,\|x\|^{2}_{\ell_{2}}:=\sum_{j=1}^{\infty}|x_{j}|^{2}<\infty\right\}.$
A classical result in this case is the following. Let $20<p<$ 2, and consider the unit ball in $\ell_{p}$,
$K=U(\ell_{p}):=\left\{{\bf x}=(x_{1},x_{2},\ldots,):\,\|x\|^{p}_{\ell_{p}}:=\sum_{j=1}^{\infty}|x_{j}|^{p}\leq 1\right\}\subset\ell_{2}.$
It is known in this case that
$\sigma_{n}(U(\ell_{p}))\asymp n^{-1/p+1/2},\quad n\to\infty,$ (1.2)
[…]
Other results of the above type have been frequently obtained in the literature. To describe a general setting, let
${\bf w}:=(w_{j})_{j\in\mathbb{N}},\quad 1\leq w_{1}\leq w_{2}\leq\dots,$
be a monotonically nondecreasing sequence of positive weights. We consider the weighted $\ell_{p}$ space $\ell_{p}({\bf w})$ defined as the set of all real valued sequences ${\bf x}\in\ell_{2}$ such that
$\ell_{p}({\bf w}):=\{{\bf x}\in\ell_{2}:\,\|{\bf x}\|_{\ell_{p}({\bf w})}<\infty\},\quad 0<p\leq\infty,$
where
$\|{\bf x}\|_{\ell_{p}({\bf w})}:=\begin{cases}\left[\sum_{j=1}^{\infty}|w_{j}x_{j}|^{p}\right]^{1/p},\quad 0<p<\infty,\\ \\ \sup_{j}|w_{j}x_{j}|,\quad\quad\quad\quad p=\infty.\end{cases}$ (1.3)
We denote by $U(\ell_{p}({\bf w}))$ the unit ball of this class
$U(\ell_{p}({\bf w}))):=\{{\bf x}\in\ell_{p}({\bf w}):\,\|{\bf x}\|_{\ell_{p}({\bf w})}\leq 1\}$
and derive the rate of the error $\sigma_{n}(U_{p}({\bf w}))$, see Theorem 2.2. It can happen that $\sigma_{n}(U_{p}({\bf w}))$ is infinite for all $n$. Also, note that the class $\ell_{p}({\bf w})$ is different from the classical weighted sequence spaces since the weights in the $\ell_{p}({\bf w})$ norm are raised to a power.
We then go on to apply our result in the case of several special weights
$w_{j}=j^{\alpha}(1+\log j)^{\beta},\quad j\in\mathbb{N},$
provided $\alpha>0,\,\beta\in\mathbb{R}$, or $\alpha=0,\,\beta\geq 0$. We show in Corollary 3.1 that when $\alpha=0$, $\beta\geq 0$,
$\sigma_{n}(U(\ell_{p}({\bf w}))\asymp n^{-(1/p-1/2)}[\log n]^{-\beta},\quad n>1,\quad 0<p<2.$
In the case $\alpha=\beta=0$, this recovers Stechkin’s result (1.2). In Corollary 3.2, we consider the case $\alpha>0$, $\beta\in\mathbb{R}$, $0<p\leq\infty$ and show that
$U(\ell_{p}({\bf w})))\asymp n^{-(\alpha+1/p-1/2)}[\log n]^{-\beta},\quad n>1,\quad\alpha+1/p-1/2>0.$
We call $\ell_{p}({\bf w})$ a generalized Wiener class in analogy with the definition of Wiener spaces in Fourier analysis when $H=L_{2}([0,1])$ and $\phi_{j}$ is the Fourier basis. Our results have some overlap with the study of Wiener classes in the Fourier setting, that is, when $\cal D$ is the Fourier basis ${\cal F}$. When considering the specific case of Fourier basis, our results, which are restricted to approximation in Hilbert spaces, are valid for $L_{2}$. Several results in the literature consider the approximation of Wiener classes in $L_{q}$ when the basis is the Fourier basis ${\cal F}$. For example, the case $\alpha>1/2$, $\beta=0$, and $p=1$ has been analyzed in [1] and upper bounds for the error in $L_{\infty}$, have been obtained for the Wiener spaces with ${\cal D}$ being the $d$-dimensional Fourier basis ${\cal F}^{d}$, see Lemma 4.3(i). Recently, these results have been improved in [2], where matching up to logarithm lower and upper bounds for multidimensional Wiener spaces with ${\cal D}={\cal F}^{d}$ are given for the case $\beta=0$, $\alpha>0$, $0<p\leq 1$, see Corollary 4.3 in [2], and $\beta=0$, $\alpha>1-1/p$, $1<p\leq q$, all when the error is measured in $L_{q}$, $2\leq q\leq\infty$, see Theorem 4.5 in [2]. In particular, when the dimension $d=1$ and the error is measured in the Hilbert space norm (i.e. $q=2$), the results from [2] give the rate $\sigma_{n}(U_{p}({\bf w},{\cal F}))_{L_{2}}\asymp n^{-(\alpha+1/p-1/2)}$, provided
$\alpha>0,\quad 0<p\leq 1,\quad\hbox{or}\quad\alpha>1-1/p,\quad 1<p\leq 2.$
This latter result is a special case of our analysis.
2 Best $n$ term approximation for $U(\ell_{p}({\bf w}))$
In going forward, we assume $H=\ell_{2}$ with its canonical basis $e_{j}$, $j\in\mathbb{N}$. Before presenting our main theorem, let us introduce the decreasing rearrangement
${\bf x^{*}}=(x_{j}^{*})_{j\in\mathbb{N}}$
of the absolute values of the coordinates $x_{j}$ of a sequence ${\bf x}=(x_{j})_{j\in\mathbb{N}}$ that is an element of the sequence space ${\bf c_{0}}$ (consisting of all sequences whose elements converge to 00). Namely, we have that $x_{1}^{*}$ is the largest of the numbers $|x_{j}|$, $j\in\mathbb{N}$, then $x_{2}^{*}$ is the next largest, and so on. It follows that
$x_{1}^{*}\geq x_{2}^{*}\geq\ldots,$
and $\|{\bf x}\|_{\ell_{p}}=\|{\bf x}^{*}\|_{\ell_{p}}$ for all $0<p\leq\infty$. For each $n\geq 1$ and ${\bf x}\in\ell_{2}$, we have that
$\sigma_{n}({\bf x})=\left[\sum_{j>n}[x_{j}^{*}]^{2}\right]^{1/2},$ (2.1)
where $\sigma_{n}({\bf x})$ is the error of $n$-term approximation of ${\bf x}$ in the $\ell_{2}$ norm. In order to prove our main result, we will need the following lemma.
Lemma 2.1.
If ${\bf x}\in\ell_{p}({\bf w})$, then ${\bf x}^{*}\in\ell_{p}({\bf w})$ , $\sigma_{n}({\bf x})=\sigma_{n}({\bf x}^{*})$, and
$\|{\bf x}^{*}\|_{\ell_{p}({\bf w})}\leq\|{\bf x}\|_{\ell_{p}({\bf w})}.$ (2.2)
Proof.
It follows directly from the definitions of $\sigma_{n}$ and ${\bf x}^{*}$ that $\sigma_{n}({\bf x})=\sigma_{n}({\bf x}^{*})$. Let the sequence ${\bf x}=(x_{1},x_{2},\dots)$ be in $\ell_{p}({\bf w})$. We can assume that all $x_{j}$ are non-negative since changing the signs of its entries does not effect neither its rearrangement nor its $\ell_{p}({\bf w})$ norm. We shall first construct sequences ${\bf y}^{(m)}=(y_{1}^{(m)},y_{2}^{(m)},\dots)$, $m=0,1,\dots$, such that each ${\bf y}^{(m+1)}$ is gotten by swapping the positions of two of the entries in ${\bf y}^{(m)}$ with the indices of these entries each larger than $m$. Also, the first $m$ entries of ${\bf y}^{(m)}$ satisfy
$y_{j}^{(m)}=x_{j}^{*},\quad j=1,\dots,m.$
Indeed, we start with ${\bf y}^{0}={\bf x}$. Assuming that ${\bf y}^{(m)}$ has been defined, we let $j>m$ be the smallest index larger than $m$ such that $y^{(m)}_{j}$ is the largest of the entries $y^{(m)}_{k}$, $k>m$. We swap the entries with positions $m+1$ and $j$ to create the sequence ${\bf y}^{(m+1)}$ from ${\bf y}^{(m)}$. We have for $m=0,1,2,\ldots,$
$\|{\bf y}^{(m+1)}\|_{\ell_{p}({\bf w})}\leq\|{\bf y}^{(m)}\|_{\ell_{p}({\bf w})}\leq\ldots\leq\|{\bf y}^{(0)}\|_{\ell_{p}({\bf w})}=\|{\bf x}\|_{\ell_{p}({\bf w})},$
because the weights in ${\bf w}$ are non-decreasing. Note that for every $m$
$\sum_{j=1}^{m}[w_{j}x^{*}_{j}]^{p}=\sum_{j=1}^{m}[w_{j}y^{(m)}_{j}]^{p}\leq\|{\bf y}^{(m)}\|^{p}_{\ell_{p}({\bf w})}\leq\|{\bf x}\|^{p}_{\ell_{p}({\bf w})},$
and therefore
$\|{\bf x}^{*}\|_{\ell_{p}({\bf w})}\leq\|{\bf x}\|_{\ell_{p}({\bf w})},$
which completes the proof.
Now we are ready to determine the rate of $\sigma_{n}(U(\ell_{p}({\bf w}))$ for all $n$ and all weight sequences ${\bf w}$. We fix ${\bf w}$, $n$, and $0<p<\infty$. From the sequence ${\bf w}$, we define the numbers
$W_{m}:=[w^{p}_{1}+w^{p}_{2}+\ldots+w^{p}_{m}]^{1/p},\quad m\geq 1.$ (2.3)
Then the following theorem holds
Theorem 2.2.
For any $0<p<\infty$ and ${\bf w}$, we have
$\max_{m\geq n}\,(m-n)[W_{m}]^{-2}\leq\sigma_{n}^{2}(U(\ell_{p}({\bf w})))\leq\max_{m\geq n}\,(m-n+1)[W_{m}]^{-2},$ (2.4)
and for $p=\infty$ we have that
$\sigma_{n}^{2}(U(\ell_{\infty}({\bf w})))\asymp\sum_{j=n+1}^{\infty}w_{j}^{-2}.$
Proof.
We fix, $n,p,{\bf w}$. For every sequence ${\bf x}\in U(\ell_{p}({\bf w}))$, we consider its decreasing rearrangement ${\bf x}^{*}$, which according to Lemma 2.1 is also an element of the unit ball and has the same $n$-term approximation. We next construct a new sequence $\tilde{\bf x}$ from ${\bf x}^{*}$ by making its first $n$ entries equal to $x^{*}_{n}$ and not touching the rest of the sequence, that is,
$\tilde{x}_{j}=\begin{cases}x^{*}_{n},\quad j=1,\ldots,n,\\ x^{*}_{j},\quad j>n.\end{cases}$
Note that because the sequence ${\bf x}^{*}$ is nonincreasing and the weights are nondecreasing, we have
$\|{\bf\tilde{x}}\|_{\ell_{p}({\bf w})}\leq\|{\bf x}^{*}\|_{\ell_{p}({\bf w})}\leq 1,\quad\hbox{and}\quad\sigma_{n}({\bf\tilde{x}})=\sigma_{n}({\bf x}^{*})=\sigma_{n}({\bf x}).$
Let us denote by $b:=x^{*}_{n}$ and notice that $\sum_{j=1}^{n}w_{j}^{p}\tilde{x}_{j}^{p}=W^{p}_{n}b^{p}$. We have
$\sigma_{n}({\bf x})=\sigma_{n}({\bf\tilde{x}})=\sum_{j=n+1}^{\infty}[\tilde{x}_{j}]^{2}=\sum_{j=n+1}^{\infty}[x^{*}_{j}]^{2}.$
We are now interested in how to change the tail of $\tilde{\bf x}$, that is, how to change the $x^{*}_{j}$’s with $j>n$ to new quantities $y_{j}$, $j>n$ so that we maximize the $n$-term approximation $\sigma^{2}_{n}({\bf y})=\sum_{j=n+1}^{\infty}[y_{j}]^{2}$, under the restrictions $b=x^{*}_{n}\geq y_{n+1}\geq\cdots$ and
$\sum_{j=n+1}^{\infty}w_{j}^{p}y_{j}^{p}\leq 1-W^{p}_{n}b^{p}=:S.$ (2.5)
Notice that an investment of $w_{j}^{p}y_{j}^{p}$ towards $S$ gives a return $y_{j}^{2}$ at coordinate $j$. Since the $w_{j}$ are non-decreasing, to maximize $\sigma_{n}({\bf y})$, it is best to invest as much as we can for $j=n+1$, $j=n+2$, and so on. So, the sequence which will maximize $\sigma^{2}_{n}({\bf y})$ has to have $y_{j}=b$, $j=n+1,\dots$ until we have used up our capital $S$. In other words, given that our sequence ${\bf y}$ has first $n$ coordinates $b$, then to maximize $\sigma_{n}({\bf y})$ we should take ${\bf y}=(b,b,\dots,b,c,0,0,\ldots)\in U(\ell_{p}({\bf w}))$, where $0\leq c<b$. The membership in the unit ball of $\ell_{p}({\bf w})$ requires that
$b^{p}\sum_{j=1}^{m}w_{j}^{p}+c^{p}w_{m+1}^{p}\leq 1\quad\Rightarrow\quad b^{p}[W_{m}]^{p}\leq 1$
and
$\sigma^{2}_{n}({\bf y})=(m-n)b^{2}+c^{2}<(m-n+1)b^{2}\leq(m-n+1)[W_{m}]^{-2}.$
Therefore, we have
$\sigma^{2}_{n}(U(\ell_{p}({\bf w})))\leq\sup_{m\geq n}(m-n+1)W_{m}^{-2}.$
Next, consider the special sequence ${\bf s}^{(m)}$ with entries ${\bf s}_{j}^{(m)}$ given by
${\bf s}_{j}^{(m)}=\begin{cases}W_{m}^{-1},\quad j=1,\ldots,m,\\ 0,\,\quad\quad j>m.\end{cases}$
Clearly ${\bf s}^{(m)}\in U(\ell_{p}({\bf w}))$ for all $m$ and
$\sigma^{2}_{n}({\bf s}^{(m)})=(m-n)[W_{m}]^{-2},\quad m>n.$
Such sequences provides the lower bound in the case $1<p<\infty$, and thus (2.4) is proven.
In the case $p=\infty$, we have that the entries $x_{j}$ of a sequence ${\bf x}\in\ell_{\infty}({\bf w})$ satisfy
$|x_{j}|\leq\|{\bf x}\|_{\ell_{\infty}({\bf w})}w_{j}^{-1}\quad\Rightarrow\quad\sigma^{2}_{n}({\bf x})\leq\left[\sum_{j=n+1}^{\infty}w_{j}^{-2}\right]\|{\bf x}\|_{\ell_{\infty}({\bf w})}.$
On the other hand, for the sequence
${\bf w}^{-1}:=(w_{1}^{-1},w_{2}^{-1},\ldots)\in U(\ell_{\infty}({\bf w})),$
we have that $\sigma^{2}_{n}({\bf w}^{-1})=\sum_{j=n+1}^{\infty}w_{j}^{-2}$, and the proof is completed.
3 Special cases of sequence spaces
In this section, we discuss several special cases of sequences ${\bf w}$ that are used in the definition of the classical Wiener spaces, see [2, 3] and the references therein.
Corollary 3.1.
Consider the classes $\ell_{p}({\bf w})$ with $w_{j}:=(1+\log j)^{\beta}$, $\beta\geq 0$, $20<p<$ 2. Then we have
$\sigma_{n}(U(\ell_{p}({\bf w})))\asymp m^{-(1/p-1/2)}[\log m]^{-\beta}.$
Proof.
Let us start by calculating the $W_{m}$,
$W^{p}_{m}=\sum_{j=1}^{m}w_{j}^{p}=\sum_{j=1}^{m}(1+\log j)^{\beta p}\geq\frac{m}{2}(1+\log(m/2))^{\beta p}.$
But we also have $W_{m}^{p}\leq m(1+\log m)^{\beta p}$, so for $m$ sufficiently big we get
$W_{m}\asymp m^{1/p}[\log m]^{\beta}.$ (3.1)
Using this in Theorem 2.2 gives that
$\sigma_{n}(U(\ell_{p}({\bf w})))\asymp n^{1/2-1/p}[\log n]^{-\beta},$
provided $20<p<$ 2.
Corollary 3.2.
Consider the classes $\ell_{p}({\bf w})$, $0<p\leq\infty$, with
$w_{j}:=\max_{1\leq i\leq j}\,i^{\alpha}\log(i+1)^{\beta},\quad j\in\mathbb{N}.$
with $\alpha>0$, $\beta\in\mathbb{R}$. Then we have
$\sigma_{n}(U(\ell_{p}({\bf w})))\asymp n^{-(\alpha+1/p-1/2)}[\log(n+1)]^{-\beta},$
provided $\alpha+1/p-1/2>0$.
Proof.
Case $0<p<\infty$: Let us observe that for any $\delta>0$ and $\gamma\in\mathbb{R}$, the function $\varphi(x):=x^{\delta}\log(x+1)^{\gamma}$, is an increasing function on $[c,\infty)$, if $c=c(\delta,\gamma)$ is sufficiently large. Therefore, we have that
$W_{m}^{p}=\sum_{j=1}^{m}j^{\alpha p}\log(j+1)^{\beta p}\asymp m^{\alpha p+1}\log(m+1)^{\beta p},\quad m\geq M,$
provided $M$ is sufficiently large. This gives
$W_{m}\asymp m^{\alpha+1/p}\log(m+1)^{\beta},\quad m\geq M.$ (3.2)
Theorem 2.2 now gives that
$\sigma_{n}(U(\ell_{p}({\bf w})))\asymp n^{1/2-\alpha-1/p}[\log(n+1)]^{-\beta},$
for $n$ sufficiently large as desired.
Case $p=\infty$: In this case we have the restriction that $12\alpha>$ 1. According to Theorem 2.2, we have that
$\sigma^{2}_{n}(U(\ell_{\infty}({\bf w})))\asymp\sum_{j=n+1}^{\infty}j^{-2\alpha}[\log(j+1)]^{-2\beta}\asymp n^{-2\alpha+1}\log(n+1)^{-2\beta}.$
Remark 3.3.
Note that when $\beta=0$, the ranges of $\alpha$ and $p$ in Corollary 3.2 are
$\alpha>0,\quad 0<p\leq 2\quad\hbox{or}\quad\alpha>1/2-1/p>0,\quad 2\leq p\leq\infty.$
Acknowledgment: This work was supported by the NSF Grant DMS 2134077 and the ONR Contract N00014-20-1-278.